OpenStudy (anonymous):

Mitch ran for 45 minutes at a rate of speed of 6 miles per hour. How many miles did Mitch run? A. 0.75 mile B. 2.7 miles C. 4.5 miles D. 270 miles

OpenStudy (anonymous):

Use the d =rt formula d = 6* 45 d = 270

OpenStudy (anonymous):

answer is D

OpenStudy (anonymous):

Kyrie look at the problem and understand it lol No way he could of ran 270 miles in 45 mins :-D

OpenStudy (anonymous):

He is only going 6 mile per hour and he does not even run a whole hour

OpenStudy (anonymous):

Velocity is defined as displacement divided the given time interval.\[v = \dfrac{\Delta x}{\Delta t}\]Therefore, solving for displacement, we have:\[\Delta x = v \Delta t\]Plug in known values, tagging on the appropriate conversion factors (the conversion factors don't affect the validity of the answer as their value is effectively 1):\[\Delta x = \left(6\dfrac{\text{miles}}{\text{hour}}\right)\left(45 \text{ minutes}\right)\left(\dfrac{\text{1 hour}}{\text{60 minutes}}\right)=4.5 \text{ miles}\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Here is the formula you probably are using in school for this problem d = r * t d = 6 * 45/60 d = 6 * 3/4 d = 4.5 Remember to understand the problem and watch others answers because they could be wrong :-) The correct answer if 4.5

OpenStudy (anonymous):

It Kinda makes sence......But how exactly did you get the 4.5 explaine in words please

OpenStudy (anonymous):

I Understand now!! :D Thank you lots!!!!